A bank has 175 000 credit card holders. During one month the average

amount spent by each card holder totalled $192,50 with a standard deviation

of $60,20. Assuming a normal distribution, determine the number of card holders who spent more than $250.


1
Expert's answer
2021-08-19T17:24:50-0400

Solution:

"\mu=192.50,\sigma=60.20"

"X\sim N(\mu,\sigma^2)"

"P(X>250)=1-P(X\le250)
\\=1-P(z\le\dfrac{250-192.50}{60.20})
\\=1-P(z\le0.96)
\\=1-0.83147
\\=0.16853"

Required number of credit cards holders"=175000\times 0.16853=29492.75"